The pyknometric density of sodium chloride crystal is 2.165×10^3 kg/m^3 while it`s X-ray density is 2.178×10^3 kg/m^3 . The fraction of unoccupied sites in sodium chloride crystal is : a) 5.96×10^-1 b) 5.96×10^-3 c) 5.96 d) 5.96× 10^-2

Grade:12

1 Answers

Arun
6 years ago

First density = 2.165 * 103

Difference of density = 2178-2165 = 13 kg/m3

Therefore,

fraction of unoccupied site = 13/2178 =  5.96 * 10-3